### 09 Apr 2026 v1

# Thermal Time and Irreversibility from Non-Commuting Observables in Accelerated Quantum Systems

Abstract
We investigate the operational meaning of temporal ordering in relativistic quantum field theory using localized detector models. The presence of a time parameter in the dynamical description does not by itself guarantee that different sequences of operations correspond to physically distinguishable processes. We show that such distinguishability arises when two conditions are simultaneously satisfied: the underlying quantum state exhibits Kubo–Martin–Schwinger (KMS) structure, and the detector couples through non-commuting observables.
We analyze uniformly accelerated two-level detectors interacting with a quantum field in the Minkowski vacuum. The restriction of the vacuum to the detector trajectory induces a thermal response characterized by the Unruh temperature and, equivalently, by a Tolman redshift profile along the trajectory. For detectors coupled sequentially through distinct internal observables, the reduced detector state depends on the ordering of the interactions. This dependence appears already at second order in perturbation theory and is controlled by the same parameter that governs the KMS condition.
The ordering asymmetry is first detected at the level of the reduced density matrices, which differ for distinct interaction protocols. Its informational content is then quantified using quantum relative entropy. In a minimal two-level model, the relevant states form a family of non-commuting Gibbs states with identical spectra and different generators. The relative entropy between these states admits a closed-form expression and depends only on the dimensionless parameter set by the local temperature and the detector energy scale.
We further analyze the corresponding information geometry by comparing the Bogoliubov–Kubo–Mori and Bures metrics. Their difference in non-commuting directions provides a quantitative distinction between irreversible entropic cost and operational distinguishability. Both quantities are governed by the same thermal scale fixed by acceleration.
These results provide an operational probe of thermal time as the modular flow generated by $K=-\log\rho$, with normalization set by the local temperature. Irreversibility arises from the mismatch between inequivalent modular structures associated with non-commuting observables, and is quantified by relative entropy. In this framework, temporal ordering becomes physically meaningful when it can be encoded in the state of a system and extracted through non-commuting measurements.
Independent researcher
marcello@gravity.phys.nagoya-u.ac.jp
Keywords thermal time $\cdot$ Unruh effect $\cdot$ KMS states $\cdot$ quantum information geometry $\cdot$ irreversibility
1 Introduction
The emergence of irreversible behavior from time-reversal invariant dynamics remains a central issue in the foundations of statistical mechanics and quantum theory. Traditional explanations rely on coarse-graining, special initial conditions, or typicality arguments, but these approaches do not directly address when a temporal ordering parameter acquires observable significance at the level of microscopic processes.
In relativistic quantum field theory, the question becomes sharper. The notion of time depends on the choice of observer, and even the classification of states as thermal or non-thermal may vary between observers. A key example is the Unruh effect, according to which a uniformly accelerated observer perceives the Minkowski vacuum as a thermal state [1, 2]. This phenomenon is naturally formulated using the Kubo–Martin–Schwinger (KMS) condition, which characterizes equilibrium states in quantum statistical mechanics [4, 5]. A detailed review is given in [3].
From an algebraic perspective, the Unruh effect is closely related to the Bisognano–Wichmann theorem, which identifies the modular flow of wedge algebras with Lorentz boosts [6, 7]. This result provides a concrete realization of modular structure in quantum field theory and underlies the thermal time hypothesis proposed by Connes and Rovelli [8]. In that framework, a state determines a canonical flow, suggesting that temporal structure is state-dependent rather than fundamental.
Detector models provide an operational framework for probing these ideas. The Unruh–DeWitt detector [9, 10] describes a localized quantum system interacting with a field along a worldline. Its response is determined by the field’s correlation functions evaluated along that trajectory. The theory of such detectors has been developed in detail, including careful treatments of switching functions and response rates [11, 12, 13].
In standard analyses, detectors couple to the field through a single observable, and the focus is on excitation probabilities and thermal spectra. This approach does not address how the internal algebra of detector observables affects the extraction of information from the field. In particular, it does not consider whether different sequences of interactions can be distinguished at the level of the detector state.
The present work focuses on this question. We consider detectors that couple sequentially through distinct internal observables that do not commute. This introduces a potential dependence on the ordering of interactions. The central result is that such ordering dependence becomes physically meaningful when the field state satisfies a KMS condition. In this case, the correlation functions distinguish forward and reversed sequences through a thermally weighted structure.
This effect admits a natural interpretation in terms of quantum information geometry. When restricted to commuting deformations of a Gibbs state, monotone metrics such as the Bures and Bogoliubov–Kubo–Mori (BKM) metrics coincide. For non-commuting deformations, they differ. The Bures metric captures operational distinguishability, while the BKM metric arises from the second-order expansion of relative entropy [14]. This distinction allows one to quantify the difference between distinguishability and irreversible entropy in a precise manner.
In the accelerated setting, the relevant dimensionless parameter is fixed by the Unruh temperature and depends on the detector’s position in Rindler space. For a minimal two-level system, the relative entropy between detector states associated with different internal observables can be computed exactly. This yields an explicit relation between acceleration, distinguishability, and irreversible entropy.
The analysis establishes a connection between three structures: KMS thermality, non-commutativity of observables, and relative entropy. Together, these determine when temporal ordering becomes observable in detector-based measurements.
The paper is organized as follows. Section 2 introduces the algebraic and detector-theoretic framework. Section 3 defines the detector model with non-commuting couplings. Section 4 analyzes the ordering dependence of the reduced detector state in accelerated motion. Section 5 provides the information-geometric interpretation in terms of relative entropy and modular structure. Section 6 summarizes the results and discusses possible extensions.
2 Algebraic and physical framework
We collect the elements required for the analysis, focusing on the structures that directly enter the detector dynamics.
Let $\mathcal{A}$ be the algebra of observables of a quantum field and let $\omega$ be a state on $\mathcal{A}$. For a scalar field $\phi$, the two-point correlation function is given by
$W(x,x^{\prime})=\omega(\phi(x)\phi(x^{\prime})).$
When evaluated along a worldline $x(\tau)$, this defines a function $W(\tau,\tau^{\prime})$ that governs the response of localized detectors. In perturbation theory, detector transition probabilities and reduced states depend only on this two-point function [10, 11, 13].
A state is said to satisfy the KMS condition at inverse temperature $\beta$ with respect to a flow parameter $\tau$ if correlation functions admit an analytic continuation and satisfy
$\omega(A(\tau)B(\tau^{\prime}))=\omega(B(\tau^{\prime})A(\tau+i\beta))$
for suitable observables $A$ and $B$ [4, 5]. This condition provides a representation-independent characterization of thermal equilibrium.
A physically relevant realization arises for uniformly accelerated observers. For a trajectory with constant proper acceleration $a$, the restriction of the Minkowski vacuum to the corresponding worldline satisfies the KMS condition with inverse temperature
$\beta=\frac{2\pi}{a}.$
This is the Unruh effect [1, 2], which can be understood algebraically through the Bisognano–Wichmann theorem [6, 7].
We model the detector as a two-level system interacting with the field through an Unruh–DeWitt coupling. The interaction Hamiltonian in the interaction picture is
$H_{I}(\tau)=\lambda\,\chi(\tau)\,\mu(\tau)\,\phi(x(\tau)),$
where $\lambda$ is a coupling constant, $\chi(\tau)$ is a switching function, and $\mu(\tau)$ acts on the detector Hilbert space. The corresponding evolution operator is
$U=\mathcal{T}\exp\left(-i\int d\tau\,H_{I}(\tau)\right),$
with $\mathcal{T}$ denoting time ordering. For an initial product state $\rho_{D}\otimes\rho_{\phi}$, the reduced detector state after the interaction is
$\rho_{D}^{\prime}=\mathrm{Tr}_{\phi}\left(U(\rho_{D}\otimes\rho_{\phi})U^{\dagger}\right).$
Expanding perturbatively, the leading contributions depend on integrals of the Wightman function along the trajectory [10, 11, 13].
The framework described above provides the connection between field correlations and detector states. In the following sections, this structure is extended by allowing the detector to couple through multiple internal observables, thereby introducing non-commutativity at the level of the detector algebra.
3 Detector model with non-commuting couplings
We now introduce the detector setup used to probe the interplay between non-commutativity and the correlation structure of the quantum field.
We consider a two-level detector with Hilbert space $\mathcal{H}_{D}\simeq\mathbb{C}^{2}$, moving along a prescribed worldline $x(\tau)$. The detector interacts with a scalar field through couplings that act on its internal degrees of freedom. In contrast with the standard formulation, we allow the detector to couple through distinct internal observables.
Specifically, we introduce two Hermitian operators
$\mu_{x}=\sigma_{x},\qquad\mu_{y}=\sigma_{y},$
where $\sigma_{x}$ and $\sigma_{y}$ are Pauli matrices satisfying $\sigma_{x},\sigma_{y}]=2i\sigma_{z}$. These operators define two independent interaction channels.
We consider interaction Hamiltonians of the form
$H_{x}(\tau)=\lambda\,\chi_{x}(\tau)\,\sigma_{x}\,\phi(x(\tau)),\qquad H_{y}(\tau)=\lambda\,\chi_{y}(\tau)\,\sigma_{y}\,\phi(x(\tau)),$
where $\chi_{x}(\tau)$ and $\chi_{y}(\tau)$ are smooth switching functions with compact, non-overlapping support. This ensures that the detector interacts with the field through $\sigma_{x}$ and $\sigma_{y}$ in a well-defined sequence. Similar perturbative constructions with localized switching are standard in detector theory [11, 12, 13].
Because the supports do not overlap, the time-ordered evolution operator factorizes as
$U_{x\to y}=U_{y}U_{x},\qquad U_{y\to x}=U_{x}U_{y},$
where
$U_{x}=\mathcal{T}\exp\left(-i\int d\tau\,H_{x}(\tau)\right),\qquad U_{y}=\mathcal{T}\exp\left(-i\int d\tau\,H_{y}(\tau)\right).$
We compare the reduced detector states obtained from the two protocols. For an initial state $\rho_{D}\otimes\rho_{\phi}$, we define
$\rho_{x\to y}=\mathrm{Tr}_{\phi}\left(U_{y}U_{x}(\rho_{D}\otimes\rho_{\phi})U_{x}^{\dagger}U_{y}^{\dagger}\right),$
$\rho_{y\to x}=\mathrm{Tr}_{\phi}\left(U_{x}U_{y}(\rho_{D}\otimes\rho_{\phi})U_{y}^{\dagger}U_{x}^{\dagger}\right).$
The difference between these two states defines the ordering-dependent contribution to the detector dynamics,
$\Delta\rho_{D}=\rho_{x\to y}-\rho_{y\to x}.$
A nonvanishing $\Delta\rho_{D}$ indicates that the two interaction protocols lead to distinct physical outcomes. The ordering of the interactions is thus encoded in the reduced detector state.
To evaluate this quantity, we expand the evolution operators perturbatively in the coupling $\lambda$ and isolate the cross terms associated with the two interaction channels. The ordering dependence arises at second order and can be expressed in terms of the commutator of the integrated interaction operators.
A detailed derivation is provided in Appendix B. The resulting ordering-dependent contribution takes the form
$\Delta\rho_{D}=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,G^{(1)}(\tau,\tau^{\prime})\,[\sigma_{z},\rho_{D}]+O(\lambda^{3}),$
where
$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
is the Hadamard function along the detector trajectory.
This expression shows that the ordering asymmetry is governed by the interplay between the non-commutativity of the detector couplings and the symmetrized correlation structure of the field.
This result makes the structure of the asymmetry explicit. The dependence on the ordering of the interactions is controlled by two ingredients: the non-commutativity of the detector couplings, encoded in $\sigma_{x},\sigma_{y}]$, and the state-dependent symmetrized field correlation function $G^{(1)}$. If either ingredient is absent, the ordering-dependent contribution vanishes.
The quantity $\Delta\rho_{D}$ establishes that the ordering of non-commuting interactions is encoded in the reduced detector state. At this stage, the result is purely dynamical: it shows that different protocols lead to different states, but does not yet quantify their distinguishability or thermodynamic significance. This will be addressed in terms of relative entropy and information geometry in the following sections.
4 Accelerated motion and KMS-induced asymmetry
We now evaluate the ordering-dependent contribution obtained in Section 3 for uniformly accelerated motion. The relevant quantity is the Hadamard function
$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau),$
whose pullback to the accelerated trajectory inherits a thermal structure from the KMS property of the vacuum state restricted to the Rindler wedge.
Let the detector follow a trajectory with constant proper acceleration $a$. The restriction of the Minkowski vacuum to this trajectory defines a KMS state at inverse temperature
$\beta=\frac{2\pi}{a}.$
Equivalently, for the family of coaccelerated Rindler observers one may write the local Tolman profile as
$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
and, for the standard Rindler metric, this yields the local inverse temperature
$\beta(\xi)=2\pi\xi$
in natural units. The KMS condition implies
$W(\tau,\tau^{\prime})=W(\tau^{\prime},\tau+i\beta),$
together with the usual analyticity properties [4, 5, 3].
For a stationary trajectory, the correlation functions depend only on the time difference, so it is convenient to introduce the Fourier transforms
$W(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{W}(\omega)e^{-i\omega(\tau-\tau^{\prime})},$
$G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)e^{-i\omega(\tau-\tau^{\prime})}.$
The KMS condition implies the detailed-balance relation
$\widetilde{W}(-\omega)=e^{-\beta\omega}\widetilde{W}(\omega),$
and therefore
$\widetilde{G}^{(1)}(\omega)=\widetilde{W}(\omega)+\widetilde{W}(-\omega)=\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega).$
If one introduces the spectral function
$\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega),$
this may also be written as
$\widetilde{G}^{(1)}(\omega)=\coth\!\left(\frac{\beta\omega}{2}\right)\widetilde{\Delta}(\omega).$
This is the thermal weighting relevant to the ordering asymmetry derived in Section 3.
Substituting the Fourier representation of $G^{(1)}$ into the exact second-order expression for $\Delta\rho_{D}$, one obtains
$\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}),$
where
$\widetilde{\chi}_{j}(\omega)=\int_{-\infty}^{\infty}d\tau\,\chi_{j}(\tau)e^{i\omega\tau}$
denotes the Fourier transform of the switching function.
Using the KMS relation, this becomes
$\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}).$
The ordering asymmetry is therefore controlled by the same local inverse temperature that determines the Unruh response and, equivalently, by the Tolman redshift profile along the accelerated trajectory.
This formula shows that the role of the KMS condition is not merely to make the detector thermal in the usual excitation-rate sense. It fixes a universal relation between the positive- and negative-frequency components of the field correlations and hence a universal thermal weighting of the symmetrized correlator that governs the ordering dependence of the reduced state.
The result also clarifies the inertial case. Along an inertial trajectory in the Minkowski vacuum, the Wightman function does not satisfy a KMS condition with respect to the detector proper time. The Hadamard function is still nonvanishing, so ordering-dependent contributions may arise from the combination of non-commutativity and field correlations. However, there is no universal detailed-balance relation and hence no thermal factor that organizes the asymmetry. Uniform acceleration singles out a regime in which the ordering dependence is governed by a well-defined local temperature.
The quantity $\Delta\rho_{D}$ therefore detects a genuine dynamical asymmetry between the two protocols, while the KMS condition endows this asymmetry with a universal thermal structure. In the next section, we turn to the question of how this state difference should be quantified informationally and how it is related to modular flow and thermal time.
5 Relative entropy, information geometry, and thermal time
The ordering asymmetry derived in the previous section admits a quantitative interpretation in terms of relative entropy and information geometry. This section has two purposes. The first is to identify the appropriate state-space quantities that measure the asymmetry detected by $\Delta\rho_{D}$. The second is to make explicit the meaning of thermal time in the present setting.
We consider the effective equilibrium states associated with detectors in a uniformly accelerated setting. The relevant temperature is not a fixed external parameter but the local temperature associated with the family of coaccelerated observers. In Rindler space this local temperature satisfies the Tolman law
$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
and for the uniformly accelerated trajectory considered here this reduces to the Unruh profile
$\beta(\xi)=2\pi\xi.$
The thermal state seen by the detector is therefore position-dependent even though the ambient state is the Minkowski vacuum.
We focus on the minimal case in which the detector couples through two distinct internal observables. Let
$H_{x}=\Delta\sigma_{x},\qquad H_{y}=\Delta\sigma_{y},$
where $\Delta$ sets the detector energy scale. The corresponding local Gibbs states are
$\rho_{x}=\frac{e^{-\beta H_{x}}}{Z},\qquad\rho_{y}=\frac{e^{-\beta H_{y}}}{Z},$
with
$Z=\mathrm{Tr}(e^{-\beta H_{x}})=\mathrm{Tr}(e^{-\beta H_{y}}).$
Introducing the dimensionless parameter
$s=\beta\Delta,$
these states take the explicit form
$\rho_{x}=\frac{e^{-s\sigma_{x}}}{2\cosh s},\qquad\rho_{y}=\frac{e^{-s\sigma_{y}}}{2\cosh s}.$
The two states have identical spectra but different generators. This is the simplest non-commuting Gibbs family relevant to the detector problem. It is also the point at which thermal time becomes nontrivial. For each Gibbs state, the modular generator is
$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
Up to the additive scalar term $\log Z$, which does not affect the modular automorphism group, the two generators are
$K_{x}\sim\beta H_{x},\qquad K_{y}\sim\beta H_{y}.$
Thermal time is therefore the flow generated by the local modular Hamiltonian, and in the present setting it depends explicitly on the local temperature through $\beta(\xi)$. The Tolman–Unruh profile determines how that modular scale varies along the accelerated trajectory.
The operator difference between the two detector states is not, by itself, an informational quantity. It detects that the two states are distinct, but it is not a canonical informational measure. The natural entropic measure is the quantum relative entropy
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y}-\rho_{y}\log\rho_{x}).$
This quantity measures the irreversible informational cost of identifying $\rho_{y}$ with $\rho_{x}$ and is therefore the appropriate object for discussing irreversibility.
Using the identities
$\log\rho_{x}=-\log(2\cosh s)\,I-s\sigma_{x},\qquad\log\rho_{y}=-\log(2\cosh s)\,I-s\sigma_{y},$
one obtains
$D(\rho_{y}\|\rho_{x})=s\,\mathrm{Tr}\big(\rho_{y}(\sigma_{x}-\sigma_{y})\big).$
The expectation value of Pauli operators in a Gibbs state of this form satisfies
$\mathrm{Tr}(\rho_{n}\sigma)=-\tanh s\,n,$
where $n$ is the unit vector specifying the direction of the generator. It follows that
$D(\rho_{y}\|\rho_{x})=s\tanh s.$
This expression provides a closed-form measure of the mismatch between the two Gibbs structures associated with the non-commuting detector couplings. It depends only on the dimensionless parameter $s=\beta\Delta$, which combines the detector energy scale with the local inverse temperature determined by the Unruh effect or, equivalently, by the Tolman profile along the accelerated trajectory.
It is important to distinguish this quantity from the perturbative state difference $\Delta\rho_{D}$ derived in Section 4. The latter establishes that different interaction orderings lead to different reduced detector states and therefore detects the presence of a dynamical asymmetry. However, $\Delta\rho_{D}$ does not by itself provide a measure of distinguishability or irreversibility.
The relative entropy $D(\rho_{y}\|\rho_{x})$ instead quantifies the distinguishability between the effective Gibbs states associated with the two non-commuting couplings. In this sense, it measures the informational content of the ordering asymmetry once expressed in terms of thermal structures.
We now relate this result to information geometry. Let $\rho_{\theta}$ denote the Gibbs state generated by
$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
so that $\theta$ parametrizes rotations between the two non-commuting generators. For nearby states, the relative entropy admits a quadratic expansion
$D(\rho_{\theta}\|\rho_{0})\simeq\frac{1}{2}g^{\mathrm{BKM}}_{\theta\theta}\theta^{2},$
where $g^{\mathrm{BKM}}$ is the Bogoliubov–Kubo–Mori metric. For the present family of states, one finds
$g^{\mathrm{BKM}}_{\theta\theta}=s\tanh s.$
By contrast, the Bures metric, which quantifies distinguishability through fidelity, yields
$g^{\mathrm{Bures}}_{\theta\theta}=\tanh^{2}s.$
The ratio between the two metrics is therefore
$\frac{g^{\mathrm{BKM}}_{\theta\theta}}{g^{\mathrm{Bures}}_{\theta\theta}}=\frac{s}{\tanh s}.$
This ratio measures the discrepancy between entropic cost and operational distinguishability for non-commuting deformations of the Gibbs state.
The dependence on acceleration can be made explicit by expressing $s$ in terms of the Rindler coordinate $\xi$. For a uniformly accelerated observer, the local inverse temperature is $\beta=2\pi\xi$, so that
$s=2\pi\xi\Delta.$
In the limit $\xi\to 0$, corresponding to large acceleration,
$D(\rho_{y}\|\rho_{x})\sim s^{2},\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\to 1.$
In the opposite limit $\xi\to\infty$, corresponding to small acceleration,
$D(\rho_{y}\|\rho_{x})\sim s,\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\sim s.$
Finally, we connect these results to modular structure. Each Gibbs state defines a modular generator
$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
Up to the additive scalar term $\log Z$, the corresponding modular flows are generated by $\beta H_{x}$ and $\beta H_{y}$. The normalization of these generators is fixed by the local inverse temperature $\beta(\xi)$, reflecting the Tolman relation and the Unruh effect.
The relative entropy $D(\rho_{y}\|\rho_{x})$ therefore measures the mismatch between the modular structures associated with the two non-commuting couplings. In this sense, thermal time is realized as the modular flow selected by the state, while irreversibility is quantified by the relative entropy between inequivalent Gibbs structures.
The detector dynamics derived in Section 4 provides the operational input: the KMS condition fixes the analytic structure of correlation functions, and non-commuting observables probe this structure through sequential interactions. The resulting ordering asymmetry is encoded in the detector state, while its informational content is captured by the relative entropy between the corresponding Gibbs states.
6 Discussion
We have identified a minimal mechanism through which temporal ordering becomes operationally meaningful in relativistic quantum systems. The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the detector couples to the field.
At the dynamical level, the ordering of interactions is encoded in the reduced detector state. Distinct interaction protocols lead to different final states whenever non-commutativity produces a nonvanishing contribution $\Delta\rho_{D}$. In the presence of a KMS state, this asymmetry acquires a universal thermal structure controlled by the local temperature. This establishes that temporal ordering can be reflected directly in the state of a localized quantum system.
The quantitative content of this asymmetry is captured by relative entropy. In the minimal two-level model, the relevant states form a family of Gibbs states with identical spectra and different generators. The relative entropy between them depends only on the dimensionless parameter $s=\beta\Delta$, linking the strength of the asymmetry to the local temperature scale.
The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the geometric structure underlying this result. In commuting directions the two metrics coincide, while in non-commuting directions they differ. This difference quantifies the separation between irreversible entropic cost and operational distinguishability, and is governed by the same thermal parameter.
Thermal time is realized in this framework through the modular generator $K=-\log\rho$ associated with a Gibbs state. In the accelerated setting, this generator is normalized by the local inverse temperature $\beta(\xi)$, determined by the Tolman relation and the Unruh effect. The dependence of the relative entropy and of the information-geometric quantities on $s=\beta\Delta$ makes explicit how irreversibility is controlled by this local thermal scale.
The detector model provides a concrete probe of these structures. The KMS condition fixes the analytic properties of the correlation functions, while non-commuting observables probe this structure through sequential interactions. The resulting effective Gibbs states encode the ordering dependence, and their relative entropy quantifies its irreversible content.
The framework suggests several directions for further investigation. One extension is to consider detector systems with higher-dimensional internal Hilbert spaces, where the structure of non-commuting generators is richer. Another is to analyze other spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where nontrivial temperature profiles may lead to additional effects.
The results show that temporal ordering acquires physical meaning when it is encoded in the state of a system through the interplay of thermality and non-commutativity. In this setting, irreversibility is quantified by relative entropy, and its magnitude is set by the same thermal scale that governs the underlying KMS structure.
Appendix A Relative entropy for non-commuting Gibbs states
We derive the closed-form expression for the relative entropy between two Gibbs states generated by non-commuting operators. This computation provides the entropic measure used in the main text to quantify the mismatch between the corresponding non-commuting Gibbs structures.
We consider a two-level system with density matrices
$\rho_{x}=\frac{e^{-s\sigma_{x}}}{2\cosh s},\qquad\rho_{y}=\frac{e^{-s\sigma_{y}}}{2\cosh s},$
where $s=\beta\Delta$.
Using the identity
$e^{-s\sigma_{n}}=\cosh s\,I-\sinh s\,\sigma_{n},$
one obtains the Bloch representation
$\rho_{n}=\frac{1}{2}\left(I-\tanh s\,\sigma_{n}\right).$
The logarithm of the density matrix follows directly from its Gibbs form,
$\log\rho_{n}=-\log(2\cosh s)\,I-s\sigma_{n}.$
The quantum relative entropy is defined as
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y})-\mathrm{Tr}(\rho_{y}\log\rho_{x}).$
Substituting the expressions above, one finds
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}\left[\rho_{y}\left(-s\sigma_{y}+s\sigma_{x}\right)\right].$
The expectation values of Pauli operators in the state $\rho_{y}$ are
$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0.$
It follows that
$D(\rho_{y}\|\rho_{x})=s\tanh s.$
More generally, if one interpolates between the two generators by
$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
the second-order expansion of relative entropy with respect to $\theta$ yields the Bogoliubov–Kubo–Mori metric used in Section 5.
The expression for $D(\rho_{y}\|\rho_{x})$ depends only on the dimensionless parameter $s=\beta\Delta$ and therefore encodes the dependence of the irreversible informational cost on the local temperature scale. In the context of the main text, this quantity provides the entropic measure of the mismatch between Gibbs states generated by non-commuting observables.
Appendix B Perturbative derivation of the ordering asymmetry
We derive the second-order ordering-dependent contribution to the reduced detector state for the two sequential interaction protocols introduced in Section 3.
Let
$A_{x}=\int d\tau\,H_{x}(\tau),\qquad A_{y}=\int d\tau\,H_{y}(\tau),$
and
$B_{x}=\int d\tau d\tau^{\prime}\,\mathcal{T}\big(H_{x}(\tau)H_{x}(\tau^{\prime})\big),\qquad B_{y}=\int d\tau d\tau^{\prime}\,\mathcal{T}\big(H_{y}(\tau)H_{y}(\tau^{\prime})\big).$
To second order in the coupling, the individual evolution operators are
$U_{x}\approx 1-iA_{x}-\frac{1}{2}B_{x},\qquad U_{y}\approx 1-iA_{y}-\frac{1}{2}B_{y}.$
Since the supports of the switching functions do not overlap, the two ordered protocols are
$U_{x\to y}\approx 1-i(A_{x}+A_{y})-\frac{1}{2}(B_{x}+B_{y})-A_{y}A_{x},$
$U_{y\to x}\approx 1-i(A_{x}+A_{y})-\frac{1}{2}(B_{x}+B_{y})-A_{x}A_{y}.$
For an initial product state $\rho_{D}\otimes\rho_{\phi}$, the reduced detector states are
$\rho_{x\to y}=\mathrm{Tr}_{\phi}\!\left(U_{x\to y}(\rho_{D}\otimes\rho_{\phi})U_{x\to y}^{\dagger}\right),$
$\rho_{y\to x}=\mathrm{Tr}_{\phi}\!\left(U_{y\to x}(\rho_{D}\otimes\rho_{\phi})U_{y\to x}^{\dagger}\right).$
At order $\lambda^{2}$, the first-order contributions and the quadratic sandwich term are identical for both protocols. The difference therefore arises entirely from the ordered cross terms and can be written as
$\Delta\rho_{D}=\rho_{x\to y}-\rho_{y\to x}=\mathrm{Tr}_{\phi}\!\left(\bigl[[A_{x},A_{y}],\,\rho_{D}\otimes\rho_{\phi}\bigr]\right)+O(\lambda^{3}).$
Using
$H_{x}(\tau)=\lambda\,\chi_{x}(\tau)\,\sigma_{x}\,\phi(x(\tau)),\qquad H_{y}(\tau^{\prime})=\lambda\,\chi_{y}(\tau^{\prime})\,\sigma_{y}\,\phi(x(\tau^{\prime})),$
we compute the commutator
$A_{x},A_{y}]=\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,[\sigma_{x}\phi(x(\tau)),\,\sigma_{y}\phi(x(\tau^{\prime}))].$
Using
$A\otimes B,\,C\otimes D]=[A,C]\otimes BD+CA\otimes[B,D],$
together with
$\sigma_{x},\sigma_{y}]=2i\sigma_{z},\qquad\sigma_{y}\sigma_{x}=-i\sigma_{z},$
one finds
$\sigma_{x}\phi(x(\tau)),\,\sigma_{y}\phi(x(\tau^{\prime}))]=i\sigma_{z}\,\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}.$
Therefore,
$A_{x},A_{y}]=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,\sigma_{z}\,\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}.$
Tracing over the field degrees of freedom gives
$\Delta\rho_{D}=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,G^{(1)}(\tau,\tau^{\prime})\,[\sigma_{z},\rho_{D}]+O(\lambda^{3}),$
where
$G^{(1)}(\tau,\tau^{\prime})=\omega\!\left(\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}\right)=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
is the Hadamard function.
For stationary trajectories, one may introduce the Fourier transform
$G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)e^{-i\omega(\tau-\tau^{\prime})}.$
If the state satisfies the KMS condition,
$\widetilde{W}(-\omega)=e^{-\beta\omega}\widetilde{W}(\omega),$
then
$\widetilde{G}^{(1)}(\omega)=\widetilde{W}(\omega)+\widetilde{W}(-\omega)=\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega)=\coth\!\left(\frac{\beta\omega}{2}\right)\widetilde{\Delta}(\omega),$
where $\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega)$.
This shows explicitly how the KMS condition induces a thermal weighting of the symmetrized correlator that governs the ordering asymmetry.

### 27 Apr 2026 v2

@@
-Thermal Time and Irreversibility from Non-Commuting Observables in Accelerated Quantum Systems
+KMS Structure, Detector Ordering, and Modular Irreversibility for Accelerated Quantum Systems
 Abstract
-We investigate the operational meaning of temporal ordering in relativistic quantum field theory using localized detector models.
-The presence of a time parameter in the dynamical description does not by itself guarantee that different sequences of operations correspond to physically distinguishable processes.
-We show that such distinguishability arises when two conditions are simultaneously satisfied: the underlying quantum state exhibits Kubo–Martin–Schwinger (KMS) structure, and the detector couples through non-commuting observables.
-We analyze uniformly accelerated two-level detectors interacting with a quantum field in the Minkowski vacuum.
-The restriction of the vacuum to the detector trajectory induces a thermal response characterized by the Unruh temperature and, equivalently, by a Tolman redshift profile along the trajectory.
-For detectors coupled sequentially through distinct internal observables, the reduced detector state depends on the ordering of the interactions.
-This dependence appears already at second order in perturbation theory and is controlled by the same parameter that governs the KMS condition.
-The ordering asymmetry is first detected at the level of the reduced density matrices, which differ for distinct interaction protocols.
-Its informational content is then quantified using quantum relative entropy.
-In a minimal two-level model, the relevant states form a family of non-commuting Gibbs states with identical spectra and different generators.
-The relative entropy between these states admits a closed-form expression and depends only on the dimensionless parameter set by the local temperature and the detector energy scale.
-We further analyze the corresponding information geometry by comparing the Bogoliubov–Kubo–Mori and Bures metrics.
-Their difference in non-commuting directions provides a quantitative distinction between irreversible entropic cost and operational distinguishability.
-Both quantities are governed by the same thermal scale fixed by acceleration.
-These results provide an operational probe of thermal time as the modular flow generated by $K=-\log\rho$, with normalization set by the local temperature.
-Irreversibility arises from the mismatch between inequivalent modular structures associated with non-commuting observables, and is quantified by relative entropy.
-In this framework, temporal ordering becomes physically meaningful when it can be encoded in the state of a system and extracted through non-commuting measurements.
-Independent researcher
-marcello@gravity.phys.nagoya-u.ac.jp
-Keywords thermal time $\cdot$ Unruh effect $\cdot$ KMS states $\cdot$ quantum information geometry $\cdot$ irreversibility
+We investigate when temporal ordering becomes physically meaningful in relativistic quantum field theory using localized detector models.
+A time parameter alone does not guarantee that different sequences of operations correspond to distinguishable processes.
+We show that distinguishability arises when two conditions are satisfied: the quantum state exhibits Kubo–Martin–Schwinger (KMS) structure and the detector couples through non-commuting observables.
+For uniformly accelerated detectors in the Minkowski vacuum, the restriction of the field state induces a thermal response characterized by the Unruh temperature and the Tolman redshift profile.
+Sequential couplings through distinct observables produce an ordering-dependent contribution to the reduced detector state already at second order in perturbation theory.
+This contribution is governed by the symmetrized correlation function and acquires a universal thermal weighting fixed by the KMS condition.
+The resulting asymmetry is quantified using relative entropy between effective Gibbs states associated with non-commuting generators.
+In a minimal two-level model, this yields a closed-form expression depending only on the dimensionless parameter set by the local temperature and detector energy scale.
+The comparison between Bogoliubov–Kubo–Mori and Bures metrics distinguishes irreversible entropic cost from operational distinguishability.
+These results provide an operational realization of thermal time as modular flow normalized by the local temperature.
+The direction of thermal time is identified with the orientation for which relative entropy production is non-negative, linking KMS thermality, modular structure, and the emergence of a thermodynamic arrow in accelerated quantum systems.
+Keywords Unruh effect $\cdot$ KMS states $\cdot$ detector models $\cdot$ modular theory $\cdot$ relative entropy $\cdot$ quantum information geometry
 1 Introduction
-The emergence of irreversible behavior from time-reversal invariant dynamics remains a central issue in the foundations of statistical mechanics and quantum theory.
-Traditional explanations rely on coarse-graining, special initial conditions, or typicality arguments, but these approaches do not directly address when a temporal ordering parameter acquires observable significance at the level of microscopic processes.
-In relativistic quantum field theory, the question becomes sharper.
-The notion of time depends on the choice of observer, and even the classification of states as thermal or non-thermal may vary between observers.
-A key example is the Unruh effect, according to which a uniformly accelerated observer perceives the Minkowski vacuum as a thermal state [1, 2].
-This phenomenon is naturally formulated using the Kubo–Martin–Schwinger (KMS) condition, which characterizes equilibrium states in quantum statistical mechanics [4, 5].
-A detailed review is given in [3].
-From an algebraic perspective, the Unruh effect is closely related to the Bisognano–Wichmann theorem, which identifies the modular flow of wedge algebras with Lorentz boosts [6, 7].
-This result provides a concrete realization of modular structure in quantum field theory and underlies the thermal time hypothesis proposed by Connes and Rovelli [8].
-In that framework, a state determines a canonical flow, suggesting that temporal structure is state-dependent rather than fundamental.
-Detector models provide an operational framework for probing these ideas.
-The Unruh–DeWitt detector [9, 10] describes a localized quantum system interacting with a field along a worldline.
-Its response is determined by the field’s correlation functions evaluated along that trajectory.
-The theory of such detectors has been developed in detail, including careful treatments of switching functions and response rates [11, 12, 13].
-In standard analyses, detectors couple to the field through a single observable, and the focus is on excitation probabilities and thermal spectra.
-This approach does not address how the internal algebra of detector observables affects the extraction of information from the field.
-In particular, it does not consider whether different sequences of interactions can be distinguished at the level of the detector state.
-The present work focuses on this question.
-We consider detectors that couple sequentially through distinct internal observables that do not commute.
-This introduces a potential dependence on the ordering of interactions.
-The central result is that such ordering dependence becomes physically meaningful when the field state satisfies a KMS condition.
-In this case, the correlation functions distinguish forward and reversed sequences through a thermally weighted structure.
-This effect admits a natural interpretation in terms of quantum information geometry.
-When restricted to commuting deformations of a Gibbs state, monotone metrics such as the Bures and Bogoliubov–Kubo–Mori (BKM) metrics coincide.
-For non-commuting deformations, they differ.
-The Bures metric captures operational distinguishability, while the BKM metric arises from the second-order expansion of relative entropy [14].
-This distinction allows one to quantify the difference between distinguishability and irreversible entropy in a precise manner.
-In the accelerated setting, the relevant dimensionless parameter is fixed by the Unruh temperature and depends on the detector’s position in Rindler space.
-For a minimal two-level system, the relative entropy between detector states associated with different internal observables can be computed exactly.
-This yields an explicit relation between acceleration, distinguishability, and irreversible entropy.
-The analysis establishes a connection between three structures: KMS thermality, non-commutativity of observables, and relative entropy.
-Together, these determine when temporal ordering becomes observable in detector-based measurements.
+Uniformly accelerated observers provide one of the clearest settings in which the relation between quantum field theory, thermality, and spacetime structure becomes operational.
+The Unruh effect shows that the Minkowski vacuum, when probed along a uniformly accelerated trajectory, gives rise to a thermal detector response with temperature proportional to the proper acceleration [1, 2, 3].
+This result is not merely a statement about detector excitation probabilities.
+It reflects the Kubo–Martin–Schwinger (KMS) structure of vacuum correlation functions restricted to accelerated motion, and is closely related to the modular properties of wedge algebras in relativistic quantum field theory [4, 5, 6, 7].
+The present paper investigates a question that arises naturally from this perspective.
+Given a detector moving along an accelerated trajectory, when does the time ordering of localized interactions become physically encoded in the detector state?
+The existence of a proper-time parameter in the interaction Hamiltonian is not sufficient by itself.
+If different interaction orderings lead to the same reduced state, the parameter orders the calculation but not an observable distinction.
+We therefore ask under what conditions the ordering defined by the detector proper time becomes operationally meaningful.
+Detector models provide a direct way to address this question.
+In the standard Unruh–DeWitt model, a localized quantum system couples to a quantum field along a prescribed worldline, and its response is determined by the pullback of the field two-point function to that trajectory [8, 9, 10, 11].
+Standard analyses usually focus on a single detector monopole operator and on transition probabilities or response rates.
+Here we instead focus on the internal algebra of the detector and allow the detector to couple sequentially through distinct non-commuting observables.
+This modification isolates a simple mechanism.
+Non-commuting detector observables make the order of operations potentially observable, while the KMS condition supplies a thermal relation between positive and negative frequency components of the field correlations.
+When both structures are present, reversing the order of the detector couplings changes the reduced detector state.
+The resulting difference is controlled by the same inverse temperature that appears in the Unruh effect.
+The second part of the paper connects this ordering asymmetry with the modular structure of effective detector Gibbs states.
+For a two-level detector with Hamiltonian generators proportional to different Pauli operators, the corresponding effective Gibbs states have the same eigenvalues but different modular generators.
+The relative entropy between these states gives a quantitative measure of the mismatch between the associated modular structures.
+In the minimal model this relative entropy can be computed exactly and depends only on the dimensionless parameter $s=\beta\Delta$, where $\Delta$ is the detector energy scale.
+This also clarifies the role of information geometry.
+The Bogoliubov–Kubo–Mori metric arises from the second-order expansion of relative entropy, whereas the Bures metric captures operational distinguishability through fidelity [12].
+For commuting thermal deformations these structures coincide in the expected classical limit, but for non-commuting deformations they differ.
+This difference provides a quantitative distinction between entropic irreversibility and state distinguishability in accelerated detector systems.
+The aim of the paper is therefore not to claim that time itself emerges from the detector model.
+A proper-time parameter is assumed throughout, as in standard relativistic detector theory.
+The result is more specific: KMS structure and non-commuting detector couplings determine when the ordering associated with that parameter becomes measurable in the detector state.
+This gives an operational interpretation of modular irreversibility in accelerated quantum systems.
 The paper is organized as follows.
-Section 2 introduces the algebraic and detector-theoretic framework.
-Section 3 defines the detector model with non-commuting couplings.
-Section 4 analyzes the ordering dependence of the reduced detector state in accelerated motion.
-Section 5 provides the information-geometric interpretation in terms of relative entropy and modular structure.
-Section 6 summarizes the results and discusses possible extensions.
+Section 2 reviews the detector and KMS framework needed for the analysis.
+Section 3 introduces the model with non-commuting detector couplings and identifies the ordering-dependent contribution to the reduced state.
+Section 4 specializes to uniformly accelerated motion and shows how the KMS condition produces a thermally weighted ordering asymmetry.
+Section 5 relates the result to relative entropy, the Bogoliubov–Kubo–Mori metric, and modular generators of effective Gibbs states.
+Section 6 summarizes the conclusions and discusses extensions.
 2 Algebraic and physical framework
@@
 $\rho_{y\to x}=\mathrm{Tr}_{\phi}\left(U_{x}U_{y}(\rho_{D}\otimes\rho_{\phi})U_{y}^{\dagger}U_{x}^{\dagger}\right).$
-The difference between these two states defines the ordering-dependent contribution to the detector dynamics,
+The difference between these two states captures the dependence on the ordering of the interactions,
 $\Delta\rho_{D}=\rho_{x\to y}-\rho_{y\to x}.$
-A nonvanishing $\Delta\rho_{D}$ indicates that the two interaction protocols lead to distinct physical outcomes. The ordering of the interactions is thus encoded in the reduced detector state.
 To evaluate this quantity, we expand the evolution operators perturbatively in the coupling $\lambda$ and isolate the cross terms associated with the two interaction channels.
@@
 where
-$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
+$G^{(1)}(\tau,\tau^{\prime})=\omega\!\left(\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}\right)=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
 is the Hadamard function along the detector trajectory.
 This expression shows that the ordering asymmetry is governed by the interplay between the non-commutativity of the detector couplings and the symmetrized correlation structure of the field.
-This result makes the structure of the asymmetry explicit.
-The dependence on the ordering of the interactions is controlled by two ingredients: the non-commutativity of the detector couplings, encoded in $\sigma_{x},\sigma_{y}]$, and the state-dependent symmetrized field correlation function $G^{(1)}$.
+The first ingredient is encoded in $\sigma_{x},\sigma_{y}]$, while the second is encoded in the state-dependent function $G^{(1)}$.
 If either ingredient is absent, the ordering-dependent contribution vanishes.
@@
 4 Accelerated motion and KMS-induced asymmetry
-We now evaluate the ordering-dependent contribution obtained in Section 3 for uniformly accelerated motion. The relevant quantity is the Hadamard function
-$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau),$
-whose pullback to the accelerated trajectory inherits a thermal structure from the KMS property of the vacuum state restricted to the Rindler wedge.
-Let the detector follow a trajectory with constant proper acceleration $a$. The restriction of the Minkowski vacuum to this trajectory defines a KMS state at inverse temperature
+We now evaluate the ordering-dependent contribution to the reduced detector state for uniformly accelerated motion.
+For a detector with constant proper acceleration $a$, the restriction of the Minkowski vacuum to the detector trajectory satisfies the KMS condition at inverse temperature
 $\beta=\frac{2\pi}{a}.$
-Equivalently, for the family of coaccelerated Rindler observers one may write the local Tolman profile as
-$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
-and, for the standard Rindler metric, this yields the local inverse temperature
-$\beta(\xi)=2\pi\xi$
-in natural units. The KMS condition implies
+The corresponding Wightman function satisfies
 $W(\tau,\tau^{\prime})=W(\tau^{\prime},\tau+i\beta),$
-together with the usual analyticity properties [4, 5, 3].
-For a stationary trajectory, the correlation functions depend only on the time difference, so it is convenient to introduce the Fourier transforms
-$W(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{W}(\omega)e^{-i\omega(\tau-\tau^{\prime})},$
+together with analyticity conditions [3, 4, 5].
+For stationary trajectories, the correlation functions depend only on the proper-time difference. It is therefore convenient to introduce the Fourier representation of the Hadamard function,
 $G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)\,e^{-i\omega(\tau-\tau^{\prime})}.$
-The KMS condition implies the detailed-balance relation
+The KMS condition implies the detailed-balance relation for the Wightman function,
 $\widetilde{W}(-\omega)=e^{-\beta\omega}\widetilde{W}(\omega),$
-and therefore
+from which it follows that
 $\widetilde{G}^{(1)}(\omega)=\widetilde{W}(\omega)+\widetilde{W}(-\omega)=\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega).$
-If one introduces the spectral function
+Equivalently, introducing the spectral function
 $\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega),$
-this may also be written as
+one may write
 $\widetilde{G}^{(1)}(\omega)=\coth\!\left(\frac{\beta\omega}{2}\right)\widetilde{\Delta}(\omega).$
-This is the thermal weighting relevant to the ordering asymmetry derived in Section 3.
-Substituting the Fourier representation of $G^{(1)}$ into the exact second-order expression for $\Delta\rho_{D}$, one obtains
+Substituting this into the expression for $\Delta\rho_{D}$, one obtains
 $\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}),$
 where
-$\widetilde{\chi}_{j}(\omega)=\int_{-\infty}^{\infty}d\tau\,\chi_{j}(\tau)e^{i\omega\tau}$
-denotes the Fourier transform of the switching function.
-Using the KMS relation, this becomes
-$\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}).$
-The ordering asymmetry is therefore controlled by the same local inverse temperature that determines the Unruh response and, equivalently, by the Tolman redshift profile along the accelerated trajectory.
-This formula shows that the role of the KMS condition is not merely to make the detector thermal in the usual excitation-rate sense.
-It fixes a universal relation between the positive- and negative-frequency components of the field correlations and hence a universal thermal weighting of the symmetrized correlator that governs the ordering dependence of the reduced state.
-The result also clarifies the inertial case.
-Along an inertial trajectory in the Minkowski vacuum, the Wightman function does not satisfy a KMS condition with respect to the detector proper time.
-The Hadamard function is still nonvanishing, so ordering-dependent contributions may arise from the combination of non-commutativity and field correlations.
-However, there is no universal detailed-balance relation and hence no thermal factor that organizes the asymmetry.
-Uniform acceleration singles out a regime in which the ordering dependence is governed by a well-defined local temperature.
-The quantity $\Delta\rho_{D}$ therefore detects a genuine dynamical asymmetry between the two protocols, while the KMS condition endows this asymmetry with a universal thermal structure.
-In the next section, we turn to the question of how this state difference should be quantified informationally and how it is related to modular flow and thermal time.
-5 Relative entropy, information geometry, and thermal time
-The ordering asymmetry derived in the previous section admits a quantitative interpretation in terms of relative entropy and information geometry.
-This section has two purposes.
-The first is to identify the appropriate state-space quantities that measure the asymmetry detected by $\Delta\rho_{D}$.
-The second is to make explicit the meaning of thermal time in the present setting.
-We consider the effective equilibrium states associated with detectors in a uniformly accelerated setting.
-The relevant temperature is not a fixed external parameter but the local temperature associated with the family of coaccelerated observers.
-In Rindler space this local temperature satisfies the Tolman law
-$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
-and for the uniformly accelerated trajectory considered here this reduces to the Unruh profile
-$\beta(\xi)=2\pi\xi.$
-The thermal state seen by the detector is therefore position-dependent even though the ambient state is the Minkowski vacuum.
-We focus on the minimal case in which the detector couples through two distinct internal observables. Let
+$\widetilde{\chi}_{j}(\omega)=\int d\tau\,\chi_{j}(\tau)e^{i\omega\tau}.$
+This expression shows that the ordering asymmetry is governed by the symmetrized correlation function of the field, with a universal thermal weighting fixed by the KMS condition.
+The relevant factor is not the difference between positive and negative frequencies, but their thermally weighted sum.
+The KMS condition therefore plays a structural role: it fixes the relation between the two frequency sectors and induces a universal thermal scaling of the ordering asymmetry.
+The magnitude of the effect is controlled by the same inverse temperature $\beta$ that determines the Unruh response.
+In the absence of KMS structure, no universal relation between $\widetilde{W}(\omega)$ and $\widetilde{W}(-\omega)$ exists, and the ordering dependence does not acquire a thermodynamic form.
+Uniform acceleration singles out a regime in which the asymmetry is governed by a well-defined local temperature.
+5 Relative entropy and information geometry
+We now quantify the ordering asymmetry in terms of relative entropy and information geometry.
+For a two-level detector in local equilibrium along an accelerated trajectory, the effective states induced by distinct coupling observables can be modeled as effective Gibbs states
+$\rho_{x}=\frac{e^{-\beta H_{x}}}{Z},\qquad\rho_{y}=\frac{e^{-\beta H_{y}}}{Z},$
+with generators
 $H_{x}=\Delta\sigma_{x},\qquad H_{y}=\Delta\sigma_{y},$
-where $\Delta$ sets the detector energy scale. The corresponding local Gibbs states are
-$\rho_{x}=\frac{e^{-\beta H_{x}}}{Z},\qquad\rho_{y}=\frac{e^{-\beta H_{y}}}{Z},$
-with
-$Z=\mathrm{Tr}(e^{-\beta H_{x}})=\mathrm{Tr}(e^{-\beta H_{y}}).$
-Introducing the dimensionless parameter
-$s=\beta\Delta,$
-these states take the explicit form
-$\rho_{x}=\frac{e^{-s\sigma_{x}}}{2\cosh s},\qquad\rho_{y}=\frac{e^{-s\sigma_{y}}}{2\cosh s}.$
-The two states have identical spectra but different generators.
-This is the simplest non-commuting Gibbs family relevant to the detector problem.
-It is also the point at which thermal time becomes nontrivial.
-For each Gibbs state, the modular generator is
-$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
-Up to the additive scalar term $\log Z$, which does not affect the modular automorphism group, the two generators are
-$K_{x}\sim\beta H_{x},\qquad K_{y}\sim\beta H_{y}.$
-Thermal time is therefore the flow generated by the local modular Hamiltonian, and in the present setting it depends explicitly on the local temperature through $\beta(\xi)$.
-The Tolman–Unruh profile determines how that modular scale varies along the accelerated trajectory.
-The operator difference between the two detector states is not, by itself, an informational quantity.
-It detects that the two states are distinct, but it is not a canonical informational measure.
-The natural entropic measure is the quantum relative entropy
-$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y}-\rho_{y}\log\rho_{x}).$
-This quantity measures the irreversible informational cost of identifying $\rho_{y}$ with $\rho_{x}$ and is therefore the appropriate object for discussing irreversibility.
-Using the identities
-$\log\rho_{x}=-\log(2\cosh s)\,I-s\sigma_{x},\qquad\log\rho_{y}=-\log(2\cosh s)\,I-s\sigma_{y},$
-one obtains
-$D(\rho_{y}\|\rho_{x})=s\,\mathrm{Tr}\big(\rho_{y}(\sigma_{x}-\sigma_{y})\big).$
-The expectation value of Pauli operators in a Gibbs state of this form satisfies
-$\mathrm{Tr}(\rho_{n}\sigma)=-\tanh s\,n,$
-where $n$ is the unit vector specifying the direction of the generator. It follows that
+and $Z=2\cosh(\beta\Delta)$.
+These Gibbs forms should be understood as effective parametrizations capturing the local thermal response induced by the KMS structure, rather than exact stationary states of the full detector-field dynamics.
+Defining $s=\beta\Delta$, these states take the form
+$\rho_{n}=\frac{1}{2}(I-\tanh s\,\sigma_{n}).$
+The quantum relative entropy
+$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y}-\rho_{y}\log\rho_{x})$
+can be evaluated explicitly, yielding
 $D(\rho_{y}\|\rho_{x})=s\tanh s.$
-This expression provides a closed-form measure of the mismatch between the two Gibbs structures associated with the non-commuting detector couplings.
-It depends only on the dimensionless parameter $s=\beta\Delta$, which combines the detector energy scale with the local inverse temperature determined by the Unruh effect or, equivalently, by the Tolman profile along the accelerated trajectory.
-It is important to distinguish this quantity from the perturbative state difference $\Delta\rho_{D}$ derived in Section 4.
-The latter establishes that different interaction orderings lead to different reduced detector states and therefore detects the presence of a dynamical asymmetry.
-However, $\Delta\rho_{D}$ does not by itself provide a measure of distinguishability or irreversibility.
-The relative entropy $D(\rho_{y}\|\rho_{x})$ instead quantifies the distinguishability between the effective Gibbs states associated with the two non-commuting couplings.
-In this sense, it measures the informational content of the ordering asymmetry once expressed in terms of thermal structures.
-We now relate this result to information geometry. Let $\rho_{\theta}$ denote the Gibbs state generated by
-$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
-so that $\theta$ parametrizes rotations between the two non-commuting generators. For nearby states, the relative entropy admits a quadratic expansion
+This quantity depends only on the dimensionless parameter $s$ and provides a measure of the mismatch between the two effective Gibbs structures associated with different generators.
+For nearby states, introduce a one-parameter family of states $\rho_{\theta}$ interpolating between the two effective Gibbs states, with $\theta$ parametrizing the rotation in operator space between $\sigma_{x}$ and $\sigma_{y}$.
+The relative entropy admits the quadratic expansion
 $D(\rho_{\theta}\|\rho_{0})\simeq\frac{1}{2}g^{\mathrm{BKM}}_{\theta\theta}\theta^{2},$
-where $g^{\mathrm{BKM}}$ is the Bogoliubov–Kubo–Mori metric. For the present family of states, one finds
-$g^{\mathrm{BKM}}_{\theta\theta}=s\tanh s.$
-By contrast, the Bures metric, which quantifies distinguishability through fidelity, yields
-$g^{\mathrm{Bures}}_{\theta\theta}=\tanh^{2}s.$
-The ratio between the two metrics is therefore
-$\frac{g^{\mathrm{BKM}}_{\theta\theta}}{g^{\mathrm{Bures}}_{\theta\theta}}=\frac{s}{\tanh s}.$
-This ratio measures the discrepancy between entropic cost and operational distinguishability for non-commuting deformations of the Gibbs state.
-The dependence on acceleration can be made explicit by expressing $s$ in terms of the Rindler coordinate $\xi$.
-For a uniformly accelerated observer, the local inverse temperature is $\beta=2\pi\xi$, so that
-$s=2\pi\xi\Delta.$
-In the limit $\xi\to 0$, corresponding to large acceleration,
-$D(\rho_{y}\|\rho_{x})\sim s^{2},\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\to 1.$
-In the opposite limit $\xi\to\infty$, corresponding to small acceleration,
-$D(\rho_{y}\|\rho_{x})\sim s,\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\sim s.$
-Finally, we connect these results to modular structure. Each Gibbs state defines a modular generator
-$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
-Up to the additive scalar term $\log Z$, the corresponding modular flows are generated by $\beta H_{x}$ and $\beta H_{y}$.
-The normalization of these generators is fixed by the local inverse temperature $\beta(\xi)$, reflecting the Tolman relation and the Unruh effect.
-The relative entropy $D(\rho_{y}\|\rho_{x})$ therefore measures the mismatch between the modular structures associated with the two non-commuting couplings.
-In this sense, thermal time is realized as the modular flow selected by the state, while irreversibility is quantified by the relative entropy between inequivalent Gibbs structures.
-The detector dynamics derived in Section 4 provides the operational input: the KMS condition fixes the analytic structure of correlation functions, and non-commuting observables probe this structure through sequential interactions.
-The resulting ordering asymmetry is encoded in the detector state, while its informational content is captured by the relative entropy between the corresponding Gibbs states.
+which for the present model yields
+$g^{\mathrm{BKM}}=s\tanh s.$
+By contrast, the Bures metric, derived from quantum fidelity, gives
+$g^{\mathrm{Bures}}=\tanh^{2}s.$
+The ratio
+$\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}=\frac{s}{\tanh s}$
+quantifies the separation between entropic irreversibility and operational distinguishability.
+At high temperature ($s\ll 1$), the two metrics coincide, while at low temperature ($s\gg 1$), they diverge.
+This behavior reflects the increasing role of non-commutativity in the structure of thermal states.
+The relative entropy therefore provides a quantitative measure of the information retained about the ordering of interactions.
+Since $D(\rho_{y}\|\rho_{x})>0$, the two states are distinguishable in principle.
+By standard results in quantum hypothesis testing, there exists an observable for which the expectation values in $\rho_{x}$ and $\rho_{y}$ differ.
+The ordering asymmetry is therefore not only a formal property of the reduced state, but is operationally accessible through suitable measurements on the detector.
+When $D(\rho_{y}\|\rho_{x})\neq 0$, the detector state encodes a distinction between different coupling sequences.
+Combined with the dynamical result of Section 4, this shows that KMS thermality and non-commutativity jointly determine when ordering information becomes physically accessible.
+The parameter $s=\beta\Delta$ controls both the strength of the Unruh response and the magnitude of the ordering asymmetry.
+The connection between modular structure and the thermodynamic arrow of time is most naturally expressed in terms of relative entropy.
+While the von Neumann entropy is conserved under unitary evolution, relative entropy with respect to a reference state provides a measure of irreversible deviation from equilibrium.
+In the present setting, the accelerated trajectory selects a local KMS state $\rho_{\mathrm{KMS}}$ and hence a modular generator $K=-\log\rho_{\mathrm{KMS}}$.
+The associated modular flow defines a natural notion of thermal time.
+The physically relevant orientation of this flow is singled out by the requirement that relative entropy production be non-negative.
+For the minimal non-commuting two-level model, this structure is reflected in the positivity of
+$D(\rho_{y}\|\rho_{x})=s\tanh s\geq 0,$
+which measures the mismatch between the effective Gibbs states associated with different coupling observables.
+The arrow of thermal time can therefore be interpreted as not imposed by the external proper-time parameter alone, but selected by the KMS state through the monotonicity of relative entropy.
+This provides a precise link between modular flow, thermality, and irreversibility: temporal ordering becomes physically oriented when it corresponds to increasing modular distinguishability.
 6 Discussion
 We have identified a minimal mechanism through which temporal ordering becomes operationally meaningful in relativistic quantum systems.
-The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the detector couples to the field.
-At the dynamical level, the ordering of interactions is encoded in the reduced detector state.
-Distinct interaction protocols lead to different final states whenever non-commutativity produces a nonvanishing contribution $\Delta\rho_{D}$.
-In the presence of a KMS state, this asymmetry acquires a universal thermal structure controlled by the local temperature.
-This establishes that temporal ordering can be reflected directly in the state of a localized quantum system.
-The quantitative content of this asymmetry is captured by relative entropy.
-In the minimal two-level model, the relevant states form a family of Gibbs states with identical spectra and different generators.
-The relative entropy between them depends only on the dimensionless parameter $s=\beta\Delta$, linking the strength of the asymmetry to the local temperature scale.
-The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the geometric structure underlying this result.
-In commuting directions the two metrics coincide, while in non-commuting directions they differ.
-This difference quantifies the separation between irreversible entropic cost and operational distinguishability, and is governed by the same thermal parameter.
-Thermal time is realized in this framework through the modular generator $K=-\log\rho$ associated with a Gibbs state.
-In the accelerated setting, this generator is normalized by the local inverse temperature $\beta(\xi)$, determined by the Tolman relation and the Unruh effect.
-The dependence of the relative entropy and of the information-geometric quantities on $s=\beta\Delta$ makes explicit how irreversibility is controlled by this local thermal scale.
-The detector model provides a concrete probe of these structures.
-The KMS condition fixes the analytic properties of the correlation functions, while non-commuting observables probe this structure through sequential interactions.
-The resulting effective Gibbs states encode the ordering dependence, and their relative entropy quantifies its irreversible content.
+The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the system probes the field.
+The mechanism can be summarized as follows.
+The KMS condition imposes a specific analytic structure on correlation functions, which translates into a frequency-dependent weighting of forward and backward processes.
+When the detector couples through non-commuting observables, this structure is reflected in the reduced detector state as a difference between distinct interaction protocols.
+The resulting asymmetry appears at the level of the density matrix and does not require coarse-graining or macroscopic limits.
+In the minimal two-level model, the detector states associated with different internal observables are described by effective Gibbs distributions with identical spectra and different generators.
+The relative entropy between these states provides a natural measure of their difference.
+Its explicit form shows that the same parameter controlling the Unruh response also governs the strength of the ordering asymmetry.
+The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the distinction between entropic and operational aspects of this effect.
+In commuting directions, these metrics coincide, reflecting the classical character of thermal fluctuations.
+In non-commuting directions, they differ, and this difference is controlled by the same parameter that determines the thermal response.
+The resulting structure links distinguishability, irreversibility, and local thermal scale in a unified way.
+The analysis presented here is based on perturbation theory and localized detector models.
+It therefore inherits the assumptions of that framework, including the use of switching functions and the restriction to leading nontrivial order in the coupling.
+Extending the results beyond this regime would require a non-perturbative treatment of detector dynamics or an explicit construction within algebraic quantum field theory.
 The framework suggests several directions for further investigation.
-One extension is to consider detector systems with higher-dimensional internal Hilbert spaces, where the structure of non-commuting generators is richer.
-Another is to analyze other spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where nontrivial temperature profiles may lead to additional effects.
-The results show that temporal ordering acquires physical meaning when it is encoded in the state of a system through the interplay of thermality and non-commutativity.
-In this setting, irreversibility is quantified by relative entropy, and its magnitude is set by the same thermal scale that governs the underlying KMS structure.
-Appendix A Relative entropy for non-commuting Gibbs states
-We derive the closed-form expression for the relative entropy between two Gibbs states generated by non-commuting operators.
-This computation provides the entropic measure used in the main text to quantify the mismatch between the corresponding non-commuting Gibbs structures.
+One natural extension is to consider more general detector systems, including higher-dimensional internal Hilbert spaces or continuous-variable models, where the structure of non-commuting observables is richer.
+Another direction is to study different spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where similar mechanisms may be present.
+The relation between modular structure and information geometry may also be explored further.
+In particular, the role of relative entropy as a measure of modular mismatch suggests a possible link with other contexts in which modular Hamiltonians play a central role, such as quantum field theory in curved spacetime and quantum information theory.
+The results show that KMS thermality, modular structure, and non-commutativity jointly determine when temporal ordering becomes observable and when it acquires a preferred direction.
+In this sense, the arrow of thermal time emerges as the direction along which relative entropy with respect to the KMS state increases.
+A natural extension of this work is to analyze whether similar ordering asymmetries arise in other settings where KMS states are present, such as stationary spacetimes with horizons.
+It would also be of interest to explore non-perturbative formulations of the effect and its relation to modular Hamiltonians in algebraic quantum field theory.
+Appendix A Derivation of relative entropy for non-commuting Gibbs states
+In this appendix we derive the closed-form expression for the relative entropy between the two effective Gibbs states introduced in Section 5.
 We consider a two-level system with density matrices
@@
 where $s=\beta\Delta$.
-Using the identity
+Using the identity for Pauli matrices
 $e^{-s\sigma_{n}}=\cosh s\,I-\sinh s\,\sigma_{n},$
-one obtains the Bloch representation
+one obtains
 $\rho_{n}=\frac{1}{2}\left(I-\tanh s\,\sigma_{n}\right).$
-The logarithm of the density matrix follows directly from its Gibbs form,
+The logarithm of the density matrix is
 $\log\rho_{n}=-\log(2\cosh s)\,I-s\sigma_{n}.$
-The quantum relative entropy is defined as
+The relative entropy is defined as
 $D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y})-\mathrm{Tr}(\rho_{y}\log\rho_{x}).$
-Substituting the expressions above, one finds
+Substituting the expressions above gives
 $D(\rho_{y}\|\rho_{x})=\mathrm{Tr}\left[\rho_{y}\left(-s\sigma_{y}+s\sigma_{x}\right)\right].$
-The expectation values of Pauli operators in the state $\rho_{y}$ are
-$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0.$
-It follows that
+Using
+$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0,$
+we obtain
 $D(\rho_{y}\|\rho_{x})=s\tanh s.$
-More generally, if one interpolates between the two generators by
-$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
-the second-order expansion of relative entropy with respect to $\theta$ yields the Bogoliubov–Kubo–Mori metric used in Section 5.
-The expression for $D(\rho_{y}\|\rho_{x})$ depends only on the dimensionless parameter $s=\beta\Delta$ and therefore encodes the dependence of the irreversible informational cost on the local temperature scale.
-In the context of the main text, this quantity provides the entropic measure of the mismatch between Gibbs states generated by non-commuting observables.
 Appendix B Perturbative derivation of the ordering asymmetry
@@
 is the Hadamard function.
-For stationary trajectories, one may introduce the Fourier transform
+For stationary trajectories, one may introduce the Fourier representation
 $G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)e^{-i\omega(\tau-\tau^{\prime})}.$
@@
 where $\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega)$.
-This shows explicitly how the KMS condition induces a thermal weighting of the symmetrized correlator that governs the ordering asymmetry.
+This shows explicitly that the ordering asymmetry is governed by the symmetrized correlation function of the field, with a thermal weighting fixed by the KMS condition.